OpenStudy (anonymous):

Which best describes the percent change? Original: 8 New:12

OpenStudy (unklerhaukus):

150%

OpenStudy (anonymous):

150%

OpenStudy (anonymous):

what how do u get that

OpenStudy (hoblos):

well the increase (12-8=4) which is half of 8 so the increase is 50%

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let x be the value to find 8x=12 x=12/8 x=1.5

OpenStudy (anonymous):

multiply bu 100 to get the percentage we get 150 percent